### v3

# Pitch: AI-assisted Tagging of Deepfake Audio Calls using Challenge-Response

Abstract.
The rise of AI voice-cloning technology, particularly audio Real-time Deepfakes (RTDFs), has intensified social engineering attacks by enabling real-time voice impersonation that bypasses conventional enrollment-based authentication. To address this, we propose Pitch, a robust challenge-response method to detect and tag interactive deepfake audio calls. We developed a comprehensive taxonomy of audio challenges based on the human auditory system, linguistics, and environmental factors, yielding 20 prospective challenges. These were tested against leading voice-cloning systems using a novel dataset comprising 18,600 original and 1.6 million deepfake samples from 100 users. Pitch’s prospective challenges enhanced machine detection capabilities to 88.7% AUROC score on the full unbalanced dataset, enabling us to shortlist 10 functional challenges that balance security and usability.
For human evaluation and subsequent analyses, we filtered a challenging, balanced subset. On this subset, human evaluators independently scored 72.6% accuracy, while machines achieved 87.7%. Acknowledging that call environments require higher human control, we aided call receivers in making decisions with them using machines. Our solution uses an early warning system to tag suspicious incoming calls as "Deepfake-likely." Contrary to prior findings, we discovered that integrating human intuition with machine precision offers complementary advantages. Our solution gave users maximum control and boosted detection accuracy to 84.5%. Evidenced by this jump in accuracy, Pitch demonstrated the potential for AI-assisted pre-screening in call verification processes, offering an adaptable and usable approach to combat real-time voice-cloning attacks. Code to reproduce and access data at https://github.com/mittalgovind/PITCH-Deepfakes
1. Introduction
Recent advancements in AI-generated voice cloning have blurred the line between authentic and fabricated speech (Li et al., 2023a; Betker, 2023; Li et al., 2023b; Wu et al., 2022; Wang et al., 2023a). This technological leap has particularly significant implications with the emergence of tools capable of generating convincing speech in real-time, fueling AI-generated impersonation scams. Such capabilities expose a critical vulnerability in billions of calls across contact centers, raising an urgent question: How can one ascertain the authenticity of a caller in an age where synthetic voices are indistinguishable from real ones?
Since 2019, the mainstream adoption of online interactions has made telephony social engineering attacks increasingly plausible. The four predominant types of phone scams, in increasing order of their targeted nature, are robocalls, replayed voices, automated voice chatbots, and real-time voice conversion. This work focuses on the latter, exploring the benefits of challenge-response mechanisms in mitigating the risks associated with real-time voice conversions.
Motivation. The scale of phone scams has reached alarming proportions. In 2022 alone, robocall-based phone scams in the United States soared to an astonishing 50.5 billion calls. By 2024, scammers had escalated their tactics, using robocalls to impersonate high-profile figures such as U.S. President Biden (nbc, 2024) and the Mayor of New York (APNews, [n. d.]) for voter suppression and outreach. This concerning escalation prompted repeated warnings from the FBI (Internet Crime Complaint Center (2021), IC3), the FTC (NPR, 2023), and an eventual ban enforced by the FCC (Federal Communications Commission, 2024).
A more sophisticated and targeted form of impersonation fraud leverages Real-time Deepfakes (RTDFs). Scammers utilize speech-to-speech (voice conversion or VC) or text-to-speech (TTS) generation tools to make deepfake calls that sound convincingly similar to a target’s voice. These real-time, interactive calls are more believable and, consequently, more persuasive than traditional scams.
While individual scam cases involving the impersonation of loved ones’ voices have been reported (Yorker, 2024), they can often be mitigated through simple precautions such as calling back, not being reactive, and asking personal questions. However, the challenge of differentiating imposters from genuine callers is far more acute and voluminous for contact centers of market entities such as banks, government agencies, and insurance companies. These organizations must continue serving millions of calls daily while protecting their clients’ sensitive accounts from impersonators.
Companies rely on various security protocols to detect impersonators, including speaker verification systems, liveness detection, and knowledge-based authentication. However, a targeted RTDF-based (live deepfake) impersonation can potentially bypass all such conventional protections (Guide, 2023; NYTimes, 2023; Online, 2023). The situation is even more dire for contact centers of organizations that lack the infrastructure to support customers’ enrollment in such protective protocols, such as insurance companies and caller-ID services like Google Voice.
While numerous audio deepfake detection methods exist (Yi et al., 2023), they are often brittle and primarily designed for static content. These traditional deepfake detectors base their judgment on a limited set of features and assume a non-interactive scenario. As these methods work offline, they are inefficient against deepfake calls where near real-time performance is desirable.
Approach. In this work, we depart from traditional techniques and leverage challenge-response mechanisms to pursue an alternative approach to detecting deepfake audio calls, which we name Pitch. Challenge-response mechanisms, traditionally pivotal in distinguishing bots from humans via the ubiquitous Captcha, have found new applications with the introduction of Gotcha (Mittal et al., 2024) and D-Captcha (Yasur et al., 2023) for unmasking video and audio deepfakes, respectively. These pioneering studies capitalize on the asymmetric advantage where the burden to maintain high quality in real-time, under challenging situations, is now squarely on the imposter caller. This work aims to harness the full spectrum of audio challenges during phone calls and orchestrate a collaborative dynamic between humans and AI, thereby fortifying accuracy and confidence.
We began by curating a detailed taxonomy of audio challenges, categorizing them into eight main types and 22 sub-types, including vocal distortions (e.g., whispering), waveform manipulations (e.g., high-pitch speaking), language-specific articulations (e.g., rolling-R sounds), environmental noise (e.g., clapping while speaking), and background playbacks (e.g., talking over music). We selected and evaluated 20 prospective challenges from the taxonomy.
Our investigation revolves around three research questions:
RQ1:
Do challenges enhance machine detection?
RQ2:
Can human evaluators harness these challenges to sharpen their discernment?
RQ3:
Does supporting humans with automated detectors further improve overall performance?
Evaluation. To address these questions, we collected a novel open-source dataset by engaging 100 participants across mobile and desktop interfaces to produce 18,600 original voice recordings. We subsequently generated 1.6 million samples through state-of-the-art one-shot voice-cloning technologies.
We performed a machine-only evaluation by designing a non-intrusive degradation metric for assessing speech samples on compliance, naturalness, and the preservation of word information. This metric revealed that a select group of 10 challenges dramatically boosted the deepfake detection performance (Area Under ROC or AUROC) to 88.7%, compared to 56.0% detecting deepfakes from regular speech, highlighting the benefit of using challenges to aid downstream machine detection. Notably, challenges like whispering, cupping mouth, and playback were among the most effective for machines.
Concurrently, we conducted a human-only evaluation on a harder balanced subset of the whole dataset by filtering 6,174 high-quality original and deepfake samples that were also positive matches to their corresponding targets. Humans assessed the compliance and quality of each speech sample and scored an AUROC of 87.7% and accuracy of 72.6%, with challenges bringing out discernible degradation in deepfakes.
Informed by these results and the propensity of humans to produce more false positives, we envisioned a framework where machines collaborated with humans in making decisions. This integration makes our approach scalable while keeping the outcome interpretable under human oversight. A subsequent human evaluation with top-performing challenges, using a balanced dataset, revealed that machine assistance increased human detection accuracy to 79.4%. Furthermore, machine assistance significantly boosted human confidence and rectified their errors in 45.4% of cases (when machines were correct) while causing misjudgments in 29.8% (when machines were wrong).
Taking this integration a step further, in scenarios where humans were uncertain, machines were allowed to take charge and make the final call. This blend of Human-AI contributions, at a ratio of 56:44, enhanced the overall detection accuracy to 84.5%, representing a 16.4% improvement over the human-only baseline (with machine taking full control standing at 87.7% accuracy). This approach underscores that while humans maintain their primary role in decision-making, strategic machine involvement can dramatically enhance the accuracy of detecting fake calls.
Our findings validated the instrumental role of challenge-response mechanisms in bolstering machine and human capabilities in audio deepfake detection. They showcased a synergy between human insight and AI accuracy that mitigates the threats posed by real-time synthetic speech technologies. As audio is always included during visual interactions, our results can transfer as is to video calls.
We list our contributions as:
A challenge-response exploration to detect real-time audio deepfakes, focusing on making it interpretable (using human-in-the-loop) and scalable (using automated methods), and extensible to more sophisticated challenges, without assuming any prior information on the suspected caller.
A tagging framework, Pitch, leverages the complementary streng -ths of humans and AI to detect deepfake calls and adapts based on prediction confidence.
A taxonomy of speech tasks comprising eight categories and 22 subcategories that are challenging for state-of-the-art voice converters, evaluated using ML models and extensive user studies to assess the efficacy and usability of these challenges.
An open-source high-quality, manually-verified dataset with 18,600 original and 1.6 million deepfake samples covering 20 diverse challenges, including code and user-study instruments.
2. Background
Definition. An audio deepfake refers to a digital impersonation in which an impostor mimics audio characteristics to convincingly match a specific target’s likeness. When such impersonations can be executed live with sufficient fidelity, we term them Real-time Deepfakes (RTDFs).
2.1. Real-time Deepfake Generation
Text-to-Speech. Text-to-Speech systems facilitate the conversion of written text into audible speech, enabling automated vocalization of textual content. The process encompasses several stages: textual analysis, phonemic conversion, and speech synthesis. Contemporary TTS methodologies employ various techniques, ranging from concatenative synthesis to advanced neural network-based approaches (suno-ai, 2023; tor, 2023). The primary objective of these systems is to generate naturalistic speech output that accurately conveys prosodic and intonational nuances inherent in human speech patterns.
Voice Conversion. VC aims to modify the linguistic features of a source speaker’s voice to emulate those of a target speaker, while maintaining the original semantic content. This process involves analysing the acoustic properties of both source and target vocalizations, followed by the establishment of a feature mapping and subsequent application of this transformation to synthesize modified speech. Recent advancements in the field have focused on real-time processing capabilities (Li et al., 2023b), the transfer of emotional vocal characteristics (Zhou et al., 2021), and the development of training methodologies utilizing non-parallel datasets (Wang et al., 2023b). A key challenge in VC research is the simultaneous preservation of speech quality and naturalness while achieving perceptually convincing voice transformations, which we aim to exploit in this work.
2.2. Cognitive Bias and Human Susceptibility
In the realm of deepfakes, individuals are often left to navigate a complex landscape independently, leaving them vulnerable to scammers who exploit their trust during phone calls. Two critical factors contribute to the rapid trust-building phenomenon in telephonic communications:
Auditory Perception Fallibility. Humans generally acknowledge the fallibility of their auditory perception. Mishearing is a common experience, arguably more prevalent than visual misinterpretation. We naturally integrate multi-sensory inputs into listening, such as live transcription, facial expressions, mouth movements, and conversational context, to enhance our ability to discern speech. These cues are sufficiently potent to augment and even override auditory perception, as demonstrated by the McGurk effect (McGurk and MacDonald, 1976). This effect illustrates how visual information from observing a speaker’s mouth movements can alter the perceived auditory signal, leading to misheard syllables. This phenomenon underscores our innate understanding that auditory signals can be unreliable, an effect that may be amplified during audio calls where the channel is noisy and visual cues are absent.
Technological Conditioning. The evolving technological landscape has significantly shaped our expectations of sound quality. Individuals accustomed to the limitations of early mobile telephony – including restricted bandwidth, dropped calls, background noise, and the need for frequent verbal confirmations – have been conditioned to interpret and understand communication even under suboptimal auditory conditions. Despite improvements in audio clarity, exposure to imperfect sound has ingrained a tolerance for noisy or unclear speech. This tolerance accentuates our inherent bias to trust and make sense of noisy speech.
This evolutionary and technological adaptation in auditory perception is evident in how humans rate audio quality. When deploying Mean Opinion Scores (MOS) for speech naturalness assessment, the scale typically ranges from 1 (incomprehensible) to 5 (comparable to face-to-face interaction). However, the average difference between the highest and lowest MOS scores is often only two points and skewed towards the higher end (Rosenberg and Ramabhadran, 2017). This pattern indicates a degree of leniency or an adaptive bias in our auditory judgments, underscoring our propensity to make sense of auditory inputs despite variations in quality.
The innate human tendency to trust rather than mistrust auditory inputs offers a potent avenue for exploitation in social engineering attacks employing audio deepfakes.
2.3. Problem Description
Threat Model. Our scenario involves three key parties: an impostor, the intended identity (the target), and a defender. The defender receives a sensitive call and anticipates the target as the caller. However, an impostor could employ an AI-generated speech synthesis tool to communicate with the defender while masquerading as the target. Given the sensitive nature of the information exchange and potential actions, the defender must authenticate the caller’s identity before proceeding. Practical examples of this scenario include bank contact centers and job interviews.
Defenders. The defender assumes no trust in the potential impostor or their devices. The defender requests that impostors perform specific tasks. While the nature of these requests is public knowledge, the task parameters are randomized at the time of request. Importantly, no identifiable and trusted target information, such as voice biometrics, is collected through an extensive enrollment process. Thus, the defender operates against a stronger threat, independent of helpful identity information.
Impostors. An impostor is assumed to have access to sufficient computational resources to run voice-cloning software in real time. They can discern and respond to requests made to them. Additionally, they can obtain a high-quality speech sample of the target, potentially via social media or an unsolicited phone call (e.g., pre-recorded voicemail greetings). This scenario is particularly impactful, as it requires minimal data yet can cause widespread harm.
Hypothesis. Speech communication comprises several vital elements, including phonetics, articulation, pitch, tone, rhythm, stress, voice quality, and fluency, all of which collectively create a distinct auditory experience. We hypothesize that an audio deepfake system should fail to maintain fidelity in real time while supporting all these speech elements simultaneously.
Distinct Concepts. It is important to differentiate our work from related but distinct concepts:
Robocalls are untargeted automated messages distributed en masse to a large population. In contrast, Real-Time Deepfakes (RTDFs), which are the focus of this work, are targeted and possibly curated for the particular receiver.
Liveness detection mechanisms are, in principle, ineffective against RTDFs, as the impostor generates deepfakes live in real-time. Moreover, we assert that the caller is a person who can converse, making each interaction unfold differently.
Works using adversarial perturbations to disrupt the deepfake generation process assume control over the impostor’s devices to inject them prior to generation. However, in our threat model, we do not trust the impostor or their devices.
3. Related Work
Deepfake Detection. Numerous techniques exist for detecting static audio deepfakes based on diverse artifacts, such as vocal tract estimation (Blue et al., 2022), Mel-frequency cepstral coefficients (MFCC) (Altalahin et al., 2023), non-semantic features (Li et al., 2024) and vocoder fingerprints (Deng et al., 2024).
However, we assert that Pitch is not a traditional ML-based deepfake detector; rather, it is a framework that can enhance the performance of current deepfake detectors by amplifying unnatural artifacts via carefully designed challenges.
Through challenge-response mechanisms, Pitch places the burden of proof squarely on the attacker. For instance, if a particular deepfake detector responds well to deepfake artifacts that sound atonal or robotic, Pitch can deploy a tonal challenge (such as “sounding happy”) against a suspected user to amplify this artifact.
Challenge-Response Systems. Authentication schemes utilizing challenge-response mechanisms are exemplified by the ubiquitous Captcha, which differentiates bots masquerading as humans on the internet. Audio-Captcha (Fanelle et al., 2020) extends this approach for people with visual impairments. The Gotcha system (Mittal et al., 2024) introduced a challenge-response approach for real-time video-only deepfake detection, analyzing eight video challenges. This study concluded that challenges lead to consistent and measurable degradation in detecting real-time deepfakes in muted videos, with a focus on interpretability.
D-Captcha (Yasur et al., 2023) developed a challenge-response system for audio calls, evaluating nine challenges and measuring detection performance based on realism, compliance, and identity against deepfake generators, with StarGANv2-VC (Li et al., 2021) being the most potent deepfake generator tested. Their research provides a vital proof-of-concept for a more systematic and robust approach, indicating the promising potential of challenge-response systems in detecting deepfake audio calls. D-Captcha++ (Nguyen-Le et al., 2024) is a follow-up that proposed an imperceptible adversarial attack on D-Captcha (Yasur et al., 2023) and employs projected gradient descent-based adversarial training to defend against it.
Human Evaluation of Deepfakes. Muller et al. (Müller et al., 2022) conducted a study with participants competing against an AI model (RawNet2 (Tak et al., 2021)) to identify audio deepfakes. Participants were shown the AI’s classification after listening to an audio clip and deciding its authenticity. The study compared human performance against AI in detecting deepfakes without allowing participants to revise their choices. The conclusion was that humans and AI exhibit similar strengths and weaknesses across various spoofing tasks from ASVSpoof2019 (Nautsch et al., 2021). Contrary to their findings, our work provides empirical evidence that human and AI performances are not strongly correlated, often complementing each other’s decision-making.
Novelty of Our Approach. To the best of our knowledge, Yasur et al. (Yasur et al., 2023) is the only closely related work that explores a challenge-response mechanism for detecting audio deepfakes. Our work expands upon this foundation in several key ways:
We explore a substantially broader spectrum of audio challenges and construct a comprehensive taxonomy based on the human auditory system, linguistics, and environmental factors.
We evaluate 20 distinct audio tasks, tested against a state-of-the-art voice converter. Notably, this modern converter rendered 6 out of 9 challenges from Yasur et al. ineffective, underscoring the rapid evolution of deepfake technology.
We introduce a novel human-AI collaborative perspective, integrating human oversight with machine efficiency to enhance detection accuracy and interpretability.
We contribute a high-quality speech dataset that significantly surpasses previous efforts in scale and diversity:
5x increase in the number of speakers
2x more challenge types
7.5x more original audio files
100x more deepfake samples
Inclusion of on-the-move smartphone environments
Comprehensive demographic diversity
These advancements extend the scope of challenge-response mechanisms in the context of audio deepfake detection into computer supported cooperative work and facilitates future research.
4. Methodology
4.1. Audio Challenges and their Taxonomy
We define an audio challenge as a task that is:
plausible to perform during a phone call,
capable of inducing degradation in real-time deepfakes,
supports randomization, and
optionally verifiable by humans.
Table 1 delineates our taxonomy of audio challenges, designed to stimulate variations in speech factors such as linguistics, tone of voice, background, and behavioral cues. The following describes each category and sub-category.
resources/evaluation/taxonomy_only
Lingual Challenges
Lingual challenges prompt individuals to alter their vocal characteristics in distinctive ways, exploiting the degrees of freedom in the human vocal system.
4.1.1. Vocal Distortions
These challenges manipulate mouth peripherals and vocal chord resonance:
Vocal Peripherals: Impeding the movement of teeth, lips, or tongue (e.g., speaking with a pen between teeth) restricts articulation, potentially hampering speech recognition in RTDFs (Zhang et al., 2022).
Whisper: Whispering, characterized by minimal vocal cord vibrations, creates a toneless output that may challenge deepfakes trained on normal speech.
Vocal Cavity: Modifying the vocal tract cavity (e.g., speaking while holding one’s nose) impacts sound resonance, potentially confounding RTDFs.
Waveform Challenges. These involve modifying key speech characteristics in the sound waveform:
Frequency: Varying voice pitch, from unusually low or high pitches to singing melodies, tests the waveform diversity an RTDF can handle (Chowdhury et al., 2022).
Amplitude: Modulating speaking volume offers a natural test for RTDF detection.
Temporal: Abnormal speech rates may challenge RTDFs trained on average speaking speeds.
Language and Articulation Challenges. These involve articulating uncommon language patterns:
Hesitation: Pronouncing difficult sequences or foreign words tests RTDFs’ ability to handle "guessing" phonetic patterns.
Mimicry: Users mimicking known accents may confound RTDFs trained on specific datasets.
Phonetics: Reproducing sounds not native to the speaker’s language (e.g., rolling Rs) challenges RTDFs’ adaptability.
Deception: Unexpected commands to cease speaking test reaction times, can expose RTDFs’ lack of human-like responsiveness.
Tone of Voice.
Emotion: Speaking with discernible emotions tests RTDFs’ ability to replicate nuanced expressions.
Phonology: Altering inflection, particularly at sentence endings, challenges RTDFs’ prosodic capabilities.
Non-lingual Challenges
These challenges rely on human-generated noises, environmental sounds, or unique speech patterns.
Noise Challenges Inducing sounds from the speaker or environment to create artifacts in deepfakes:
Vocal Noises: Includes coughs, whistles, and loud breathing.
Non-vocal Noises: Involves producing intentional sounds like clapping, snapping fingers, or microphone manipulation.
Environmental Noises: Incorporates background sounds like bird songs or traffic noise.
Reverb: Creates feedback loops or reverberations during the call.
Behavioral Challenges. (out-of-scope) These utilize known attributes of an individual’s speaking style or biometrics but are not explored in this work due to our assumption of having no prior identity information.
Replay Challenges
Playback. This category involves playing recordings that may disrupt deepfake creation:
Speech: Subjects speak while a recorded voice is playing, potentially creating artifacts as RTDFs struggle with multiple speakers.
Music: Speaking over music introduces non-speech audio patterns, challenging the voice conversion process.
Passive Distortions (out-of-scope): They involve manipulating the caller’s voice or playing adversarial audio, are considered out-of-scope as they assume trust in the caller’s equipment.
4.2. Challenges Dataset Curation
Ethics and Recruitment. Following our Institutional Review Board (IRB) approval, we recruited participants using Prolific. We informed all subjects about the intended use of their data for creating and evaluating deepfakes. Participants provided consent for video and audio recording during task performance and agreed to share their data for future secondary research.
We carefully considered the ethical implications of creating and using deepfakes, including potential misuse and privacy concerns. To mitigate these risks, we implemented strict data handling protocols and enforcement to access the generated deepfakes.
Study Design. Our study encompassed recordings in both desktop and smartphone environments. In the desktop setting, we explored 21 tasks, including standard speech in challenge #0, i.e. the "no-challenge" condition. For 20 of the 21 challenges, each participant read ten phonetically diverse sentences per challenge (see Table A.1). We chose ten sentences per challenge to balance comprehensive data collection with participant fatigue. We limited challenges involving coughing or whistling to one instance. Consequently, each successful desktop participant recorded 201 samples.
We designed the smartphone study to simulate a mobile user scenario with potentially limited task-performance freedom. As such, the smartphone dataset constituted a strict subset of the desktop dataset. We excluded playback challenges (#18, #19, #20), which necessitated a secondary device for background sound playback—an unrealistic assumption in a mobile context. Thus, each successful smartphone participant recorded 171 samples.
We developed the experiments using the Gorilla Experiment Builder (see Fig A.4).***We open our instrument at https://app.gorilla.sc/openmaterials/722500. The initial data collection involved 100 subjects, equally distributed between smartphone and desktop environments, yielding 18,600 original recordings. This dataset is comparable in scale to the well-known VCTK speech dataset (Yamagishi et al., 2019), which contains 110 speakers with 400 samples each (totaling 44,000 recordings), while offering significantly greater diversity. Our participant pool consisted of individuals aged 18-65, near equal gender distribution, and is multi-racial. This diverse demographic helped ensure a representative sample of speech patterns and characteristics.
Data Pre-processing. We extracted the 48kHz audio stream from the recorded samples (.webm) and down-sampled them to 16kHz lossless audio (.wav) files for downstream tasks. We employed WhisperX (Bain et al., 2023) to transcribe the down-sampled audio samples and obtain timestamps for speech segments. We then concatenated the segmented parts to clip the speech’s start and end, effectively removing silent intervals.
Manual Quality Check of Original Recordings: To ensure dataset quality, we engaged 45 human evaluators to verify the presence of required challenges in the pre-processed audio recordings. To optimize the evaluation process, we randomly sampled one audio recording per subject per challenge for compliance checking, assuming consistent performance across multiple attempts.
Three English-speaking evaluators—a native monolingual, a non-native multilingual, and a tie-breaker—verified each sample. This approach aimed to enhance inclusivity and mitigate demographic bias. We made compliance decisions using majority voting, resulting in an 84% mean compliance across 100 subjects and 89% across the 21 challenges. Fig. A.6 illustrates the subject-compliance distribution across the dataset and demographic distribution.
Choosing RTDF generators. Our analysis revealed that state-of-the-art Text-to-Speech (TTS) models such as TorToise (Betker, 2023) and Bark (suno-ai, 2023) excelled in content modulation but fell short in capturing speech nuances like emotions or prosody, significantly limiting their effectiveness against our challenges. Consequently, we shifted our focus to speech-to-speech systems, which convert source speech into the targeted speaker’s voice while preserving subtle speech elements.
Among the various voice converters available, we selected FREE-VC (Li et al., 2023b), PPG-VC (Liu et al., 2021), and StarGANv2-VC (Li et al., 2021) for our evaluation. We most thoroughly evaluate FREE-VC, as it is open-source and has any-to-any voice cloning capabilities, outperforming earlier models and matching recent advancements. A threat analysis study by Korshunov et al. (Korshunov et al., 2023) identified it as a high threat to speaker recognition systems, suggesting its potential as an imposter’s preferred choice. PPG-VC represented a prior state-of-the-art deepfake generator, while StarGANv2-VC was the subject of evaluation in previous work by Yasur et al. (Yasur et al., 2023).
Generating Deepfakes: For deepfake generation, we considered each of our 100 participants as a target, with all others serving as imposters, resulting in unique target-imposter pairs. Our challenges utilized one-sentence (s) samples of both target and imposter speech. This setup intentionally varied speech content between target and imposter to reflect realistic constraints. We converted all collected samples into deepfakes using FreeVC (Li et al., 2023b), generating a total of 1.6 million samples (99 imposters 18,600 samples, minus 241,000 failure cases).
With the challenges selected and speech dataset of original and deepfakes in place, we evaluate them next with machines, humans and subsequently their collaboration.
| Index | Challenge | Human | Compliance Check using Finetuned Wav2Vec2 | |||||
| Baseline | Val Acc. | Test Acc. | AUC | Precision | Recall | F1 | ||
| #1 | Static Mouth | 85.9 | 88.2 | 96.6 | 0.89 | 0.87 | 0.88 | |
| #2 | Cup mouth | 86.9 | 89.5 | 97.1 | 0.95 | 0.84 | 0.89 | |
| #3 | Whisper | 94.9 | 90.9 | 93.1 | 0.92 | 0.9 | 0.91 | |
| #4 | Hold nose | 43.4 | 90.0 | 92.8 | 0.94 | 0.85 | 0.90 | |
| #5 | High Pitch | 86.9 | 81.8 | 91.2 | 0.95 | 0.67 | 0.79 | |
| #6 | Low Pitch | 73.7 | 77.3 | 85.9 | 0.75 | 0.83 | 0.78 | |
| #7 | Sing | 87.9 | 90.0 | 96.8 | 0.89 | 0.91 | 0.9 | |
| #8 | Speak Loudly | 86.9 | 70.5 | 74.8 | 0.73 | 0.65 | 0.69 | |
| #9 | Speak softly | 85.9 | 79.1 | 84.7 | 0.76 | 0.85 | 0.8 | |
| #10 | Read quickly | 98.0 | 83.2 | 90.7 | 0.92 | 0.73 | 0.81 | |
| #11 | Read Slowly | 89.9 | 79.1 | 87.5 | 0.8 | 0.77 | 0.79 | |
| #13 | Accent | 72.7 | 79.5 | 85.9 | 0.81 | 0.77 | 0.79 | |
| #14 | Emotions | 89.9 | 88.6 | 95.6 | 0.98 | 0.79 | 0.87 | |
| #16 | Cough/Whistle | 90.9 | 96.7 | 99.3 | 0.82 | 0.82 | 0.82 | |
| #17 | Clap | 84.7 | 94.1 | 98.8 | 0.95 | 0.93 | 0.94 | |
| #18 | Cross-talk | 91.5 | 98.8 | 100.0 | 1.0 | 0.96 | 0.98 | |
| #19 | Instr. Playback | 95.7 | 98.1 | 99.6 | 0.98 | 0.96 | 0.97 | |
| #20 | Lyric Playback | 100.0 | 98.1 | 100.0 | 0.94 | 1.0 | 0.97 | |
| Average (across all) | 86.0 | 87.5 | 93.1 | 0.89 | 0.82 | 0.85 | ||
| Heuristic-based Compliance Check | ||||||||
| #12 | Foreign Words | 99.0 | WER compared to the specific foreign script used. | |||||
| #15 | Question | 92.9 | Transcribed sentence has a question mark | |||||
5. Evaluation
5.1. RQ1. Machine Evaluation
We evaluate the effectiveness of our framework against voice-cloning attacks using an automated authentication system. An impersonator’s goal is to mimic a target individual during a call by:
Adhering to the challenge in terms of task and content,
Maintaining the target speaker’s identity, and
Achieving a high level of realism.
Conversely, our challenges aim to disrupt these goals. To assess challenge performance, we quantify the disruption caused. However, incorporating speaker identity verification would require a reference sample, making the evaluation intrusive (needing a reference sample of the target’s voice) and thus, impractical. Therefore, we focus on three non-intrusive criteria:
These criteria were selected for their ability to comprehensively assess voice sample quality and authenticity without reference samples, maintaining the non-intrusive nature of our evaluation.
Challenge Compliance . We assess this criterion using a set of 18 binary classifiers, trained on the original speech data described in §4.2. We designed each classifier to distinguish between a specific challenge and regular speech (challenge #0). We partitioned the dataset by subject identity for 5-fold cross-validation, using 90 identities and reserving 10 for a final test set.
We based each classifier on a fine-tuned Wav2Vec2 model (Baevski et al., 2020). We chose Wav2Vec2 as the feature backbone, as it was pretrained on 50,000 hours of unlabeled speech and provided superior performance against MFCC, LFCC or Mel Spectrogram-based classifiers.
We fine-tuned each Wav2Vec2 model for 30 epochs with a batch size of 64 and a learning rate of 5e-5. During training, we monitored validation accuracy after each epoch, retaining the best-performing model. This process yielded mean validation and test accuracy of 87.5% and 83.6% 8.0%, respectively, and achieved a mean AUROC of 93.1%. For context, human evaluators rated compliance at 86.0% (provided by human evaluators to the original challenges). Thus, our models are close to ground truth accuracy.
For the remaining 2 out of 20 challenges—speaking foreign words (challenge #12) and asking questions (challenge #15)—we assessed compliance through speech transcription (see Table 2).
While we follow machine learning protocols for dataset splitting, we apply the compliance detectors to all identities. Importantly, we do not use any deepfakes for training or validation, as we are only testing compliance.
Content Compliance. We employ Word Information Lost (WIL) (Morris et al., 2004) to quantify the discrepancy between transcribed speech content and the original script. WIL was chosen over standard Word Error Rate (WER) for its emphasis on keywords and de-emphasis of fillers, better reflecting natural conversation dynamics. This metric is applied to all challenges except those involving non-verbal sounds (e.g., coughing or whistling).
Realism . We utilize the Non-Intrusive Speech Quality Assessment (NISQA) (Mittag et al., 2021) to evaluate realism. NISQA is an ML framework for speech quality prediction, which was trained and evaluated on 81 diverse datasets. It has been extensively adopted and recognized by the Speech community. NISQA predicts a 5-point mean opinion score (pMOS) for a given audio sample without requiring a reference signal, aligning with general speech research applications.
| No. | Challenge | WIL | AUC | ||||
| #0 | No Challenge | - | 8 | 11 | 11 | 5 | 56 |
| #1 | Static Mouth | 22 | 10 | 57 | 31 | 11 | 85 |
| #2 | Cup mouth | 87 | 11 | 38 | 46 | 9 | 97 |
| #3 | Whisper | 61 | 27 | 63 | 53 | 22 | 92 |
| #4 | Hold nose | 55 | 11 | 32 | 29 | 38 | 63 |
| #5 | High Pitch | 63 | 9 | 28 | 44 | 17 | 86 |
| #6 | Low Pitch | 35 | 10 | 20 | 25 | 15 | 65 |
| #7 | Sing | 18 | 10 | 32 | 21 | 9 | 70 |
| #8 | Speak Loudly | 34 | 7 | 13 | 26 | 16 | 62 |
| #9 | Speak softly | 36 | 12 | 22 | 29 | 16 | 71 |
| #10 | Read quickly | 12 | 8 | 35 | 24 | 14 | 67 |
| #11 | Read Slowly | 19 | 11 | 13 | 18 | 20 | 55 |
| #12 | Foreign Words | - | 10 | 82 | 47 | 6 | 97 |
| #13 | Accent | 25 | 9 | 26 | 21 | 18 | 57 |
| #14 | Emotions | 30 | 9 | 25 | 35 | 10 | 82 |
| #16 | Cough/Whistle | 36 | 26 | - | 38 | 34 | 54 |
| #17 | Clap | 12 | 24 | 33 | 22 | 19 | 51 |
| #18 | Crosstalk | 27 | 17 | 65 | 38 | 13 | 90 |
| #19 | Instr. Playback | 96 | 24 | 40 | 53 | 17 | 98 |
| #20 | Lyric Playback | 26 | 25 | 61 | 38 | 17 | 85 |
| Average | 37.3 | 13.6 | 35.8 | 31.7 | 15.9 | 74.4 | |
| Average (top-10) | 49.5 | 15.4 | 48.0 | 39.5 | 13.8 | 88.7 | |
Given these three measures, we define a Machine-scored degradation for an audio sample (x) and challenge (chal). A higher score indicates poorer deepfake performance, aligning with our challenge objectives. The degradation score is calculated as:
where indicates compliance with the challenge, returning 1 for compliance and 0 otherwise.
We assessed the degradation score across 1.6 million fake and 18,600 original audio samples. Due to the significant imbalance between fake and original samples in our automated evaluation dataset, we rely on AUROC rather than accuracy as our primary metric. This choice ensures a more robust evaluation given the dataset’s imbalanced nature.
Table 3 details our analysis, highlighting the degradation disparity between fake and original samples for all challenges. Fig. 2 visually demonstrates the separation in machine-scored degradation between original and fake samples.
Observations. Regular speech (no-challenge) exhibited low degradation—11% for deepfakes and 5% for originals—indicating RTDFs’ proficiency in replicating standard speech patterns.
The introduction of audio challenges led to increased degradation in both deepfakes and original recordings, with deepfakes experiencing more substantial quality deterioration. This effect was particularly pronounced in tasks involving foreign words, static mouth, and cross-talk, resulting in heavy word information loss.
Our examination of 20 prospective challenges revealed 10 to be less effective, with AUROC scores below 70% (Table 3), leading to their disqualification. These tasks included variations in accent, reading pace, amplitude, pitch, and specific actions like clapping, coughing/whistling, and nose-holding. Deepfake algorithms demonstrated notable proficiency in mimicking tasks related to regular speech patterns, showcasing the sophistication of voice-cloning systems in replicating human vocal characteristics.
Additional disqualification reasons included: (1) low compliance rate in original recordings for the hold-the-nose challenge (43%), (2) non-randomizability of the cough/whistle challenge due to lack of spoken content, and (3) false positives in the clapping challenge detector, where deepfakes transformed clap sounds into noisy peaks mimicking clap-like features in the waveform.
Deepfakes struggled to maintain compliance against top challenges in scenarios requiring lyrical playback, mouth cupping, whispering, and high-pitched voice production (Table 3). These findings indicate a critical vulnerability in current deepfake technologies. Lyrical playback, in particular, introduced measurable vocal distortions, highlighting the effectiveness of these challenges in compromising deepfake audio fidelity.
Our analysis of frequency modulation effects, comparing high-pitch and low-pitch challenges, revealed that high-pitch challenges induced more significant degradation (44%) in deepfake outputs than their low-pitch counterparts (25%). This discrepancy in machine detection efficiency points to limitations in the training datasets of deepfake models, such as the VCTK corpus, which lacks diverse high-pitch voice samples. This limitation underscores the critical role of dataset diversity in enhancing deepfake quality.
Despite the underperformance of specific tasks, we included them in our analysis for comprehensive evaluation and comparison with Yasur et al. (Yasur et al., 2023). The authors had identified six out of the nine ineffective tasks as beneficial against the StarGANv2-VC model (Li et al., 2021). However, these tasks proved less effective when tested against the more sophisticated FREE-VC model (Li et al., 2023b), suggesting that our remaining 10 challenges offer higher degradation even against advanced adversaries.
Takeaways. We have developed a comprehensive and robust machine evaluation framework, identifying 10 tasks that effectively serve as challenges. These challenges enable machines to achieve an AUROC of 88.7%, compared to 56.0% for no-challenge conditions. Our findings highlight the vital role of challenges in identifying deepfake strengths and limitations, particularly when such tasks involve complex auditory challenges beyond simple speech imitation. The analysis reveals two main vulnerabilities of RTDFs: they are restricted by training data diversity, and they do not encode the physiology and physics of human speech apparatus.
Additional Results (Table A.2). We also evaluate two popular deepfake detectors: LFCC feature extraction with LCNN classifier (Liu et al., 2022) and RawNet2 (Tak et al., 2021)/SpecRNet, scoring 75.1% and 86.0% AUC, respectively. Our machine evaluation bypasses them.
Additionally, we evaluate two any-to-many voice converters: StarGANv2-VC (Li et al., 2021) and PPG-VC (Liu et al., 2021). We trained them on our open-sourced dataset to add new identities to their target group. The original implementation of StarGANv2-VC only included 20 speaker identities. In comparison, we trained it on our 100 speaker identities, making them all seen during training. The challenges remain effective as our automated detectors scored AUC of 88.2% and 88.3% against StarGANv2 and PPG, respectively.
5.2. RQ2. Vanilla Human Evaluation
While we have established that challenges enable effective machine evaluations, it is crucial to investigate whether similar results hold for human assessments. System support for human evaluation is essential in scenarios where advanced deepfake detection tools are unavailable, interpretability is desired, or both. This subsection details our vanilla human evaluation and its subsequent extension into two machine-assisted evaluations.
Creating a Balanced Dataset. Given the impracticality of manually evaluating 1.6 million audio files, we focused on a hard balanced subset of the data.
Our full dataset comprises 9,900 unique imposter-target speaker pairs. We employed the SpeechBrain speaker recognition framework (Ravanelli et al., 2021) to compute voice matching scores () for each pair while maintaining consistent speech content between samples. Of these, 5,643 imposter-target pairs matched at a threshold of 0.25, where a given speech sample is classified as a positive match of the reference. To increase the likelihood of human failure, we further restricted the set by raising the threshold to 0.50. From this subset, we selected 147 high-quality samples per challenge (21 in total), each with a (pMOS) of (), resulting in a hard balanced dataset of 6,174 audio samples for human assessment.
Setup. Our human evaluation methodology mirrors the compliance check for original recordings outlined in §4.2. Therefore, we recruited 63 English-speaking evaluators, categorized into three groups: native monolingual, non-native multilingual, and tie-breaker. We determine the final score for each sample through majority voting across three evaluators.
Evaluators were tasked with assessing each audio sample’s compliance with the given task () and its overall quality or realism () on a 5-point scale, following the ITU P.808 standard for subjective audio quality assessments (ITU, 2018). See Fig. A.3. †††Reproducible at https://app.gorilla.sc/openmaterials/769577
We calculate the human-scored degradation for an audio sample as follows:
Here, yields 1 if compliant and 0 otherwise. This measure assigns maximum degradation to a non-compliant sample or a score reduction based on human-rated quality.
Fig. 3 illustrates human evaluation results for the top-10 challenges, while Fig. A.1 illustrates the distribution of human scores across the 20 challenges and the corresponding AUROC.
Humans vs Machines. Humans scored overall 74.7% AUROC on the 20 challenges, with strong human-human agreement (Kendall’s Coefficient ).
Human performance surpassed machines in challenges involving Speaking Loudly, Low Pitch, Cough/Whistle, and Clapping. Conversely, humans were less effective in challenges associated with Static Mouth, Foreign Words, Singing, and Emotion Display, indicating a human sensitivity gap in detecting subtleties within these challenges.
| (a) Human decision | (b) Machine verdict | (c) Chance to Update decision |
On the samples corresponding to the top-10 challenges in the balanced dataset, human evaluators achieved a mean AUROC of 87.7% in detecting deepfakes across, comparable to machine evaluations (88.7%), albeit on a difficult, high-quality dataset. However, their accuracy (not AUROC) was lower at 72.6% mean accuracy, in comparison to machines’ 87.7%. See breakdown in first column of Table 5. Interestingly, for this set, the Pearson correlation between human and machine AUROC was merely 0.12, implying a poor positive relationship and revealing the complementary strengths of humans and machines.
Also, the machine realism (NISQA) and Humans’ MOS ratings on our dataset were uncorrelated or weakly uncorrelated (Pearson Correlation ), corroborating our finding that machines and humans are sensitive to different aspects of speech. See Tab. A.2 for breakdown of coefficients.
Key Takeaway. Although the disparity between human and machine performance appears nominal, each exhibits distinct advantages and limitations. This gap becomes more pronounced when examining the accuracy on the 10 machine-validated challenges. Here, humans achieved an accuracy of 72.6% in deepfake detection, compared to 87.7% by machines. This difference raises the prospect of combining human and machine efforts to enhance overall performance in deepfake detection.
| Challenge | Vanilla Acc | Assisted Acc. | Collaborative | ||
| No Challenge | 52.2 | 66.7 | 27.2% | 88.4 | 69.7% |
| Static Mouth | 58.8 | 68.8 | 16.2% | 71.1 | 20.9% |
| Cup mouth | 84.9 | 91.2 | 7.4% | 93.9 | 10.6% |
| Whisper | 69.6 | 78.1 | 12.6% | 81.7 | 17.9% |
| Speak softly | 61.8 | 63.4 | 2.4% | 66.5 | 7.5% |
| High Pitch | 84.5 | 90.5 | 6.9% | 90.6 | 7.1% |
| Foreign Words | 58.5 | 71.3 | 21.5% | 85.4 | 46.3% |
| Emotions | 68.3 | 74.0 | 7.9% | 77.3 | 12.3% |
| Crosstalk | 80.2 | 85.1 | 6.0% | 86.3 | 7.2% |
| Instr. Playback | 89.2 | 94.6 | 5.8% | 95.1 | 6.3% |
| Lyric Playback | 90.2 | 90.7 | 0.5% | 89.1 | 1.2% |
| Average Acc. | 72.6 | 79.4 | 9.3% | 84.3 | 16.4% |
5.3. RQ3. AI Assistance & Human Collaboration
Given the complementary strengths exhibited by humans and machines, we conducted additional human evaluations to explore the potential of machine assistance in enhancing human capability to detect deepfakes. We focused this investigation on the top 10 challenges (and no-challenge) where machine performance was deemed satisfactory.
For this evaluation, we selected a subset of 92 samples from the original 147 in the hard-balanced dataset. This subsampling resulted in 2,016 deepfake and original audio samples, and allowed us to recruit a larger pool of evaluators.
Study Design. We recruited a new, gender-balanced cohort of 100 human evaluators from the Prolific platform. The recruitment process included a short orientation to familiarize participants with the challenge samples they would encounter. We used a preliminary quiz based on original challenge samples to screen participants, with 91 qualifying to serve as an evaluator.
We informed participants about the task and emphasized that the machine’s verdict had 87.7% accuracy (computed by choosing as the threshold for all challenges). Each participant was asked to listen to an audio sample and determine its authenticity ("Real" or "Fake"), as well as their confidence level in this initial decision. After making their initial judgment, we disclosed the machine’s verdict to the participant, sometimes accompanied by a brief explanation. We based the explanations on one of three criteria established in §5.1: failure to perform the specified task, mismatch in spoken text, or presence of vocal distortions. See Fig. 4.
After receiving the machine’s verdict and rationale, participants were given the option to revise their decision and confidence level or stick with their initial assessment. This method generated a total of 8,372 responses.
✓: correct, ✗: incorrect. Rightmost column: mean SD of confidence change.
| Scenario | Count | I | M | F | Conf.(%) |
| Correct Agreement | 5362 | ✓ | ✓ | ✓ | |
| Machine Corrected | 781 | ✗ | ✓ | ✓ | |
| Self Corrected | 5 | ✗ | ✗ | ✓ | |
| No Change (Initial Correct) | 510 | ✓ | ✗ | ✓ | |
| Self Misled | 27 | ✓ | ✓ | ✗ | |
| Machine Misled | 216 | ✓ | ✗ | ✗ | |
| No Change (Initial Wrong) | 939 | ✗ | ✓ | ✗ | |
| Incorrect Agreement | 532 | ✗ | ✗ | ✗ |
Machine-Assisting Humans. Table 4 (center) details the accuracy of each challenge after machine assistance. We report accuracy rather than AUROC, as the dataset is perfectly balanced. Vanilla decisions achieved a 72.6% accuracy, while assisted decisions improved to 79.4%, enhancing the vanilla accuracy by 9.3% overall. Notably, machine assistance increased vanilla human accuracy across all individual challenges. This effect is particularly evident in nuanced challenges such as Static-Mouth, Cupping-Mouth, and Whisper, and less pronounced in scenarios like No-Challenge and Foreign Words (as evaluators were English speakers).
Table 5 presents the initial judgments (I), machine-assists (M), and final decisions (F) across the challenges, breaking down the responses across eight possible scenarios and noting changes in confidence levels. When machine predictions aligned with human decisions, they increased human confidence, regardless of correctness. Conversely, when predictions diverged, they reduced human confidence. Notably, when humans initially erred and machines provided correct guidance, humans changed their decisions 45.4% of the time (781 / (781 + 939)). However, when humans were correct and machines erred, the latter misled humans 29.8% of the time (216 / (216 + 510)). This disparity highlights the beneficial role of machines in correcting human errors and their relatively lesser impact in misleading humans when they are correct.
Comparing the presence of rationale against providing no rationale, we observe no significant difference in humans accepting the decision and changing their mistakes (42.4% vs. 43.0%).
Human-AI Collaboration. The machine-assistance framework positioned machines as a warning signal, preventing them from making unilateral final decisions. This approach necessitates convincing humans to correct their initial judgments based on machine input, which could be cumbersome. To bypass this, we introduced a mechanism allowing the machine to override human decisions when it exhibited higher confidence (not correctness of decision).
We computed the machine confidence for each audio sample’s decision using the machine scores from §5.1. The machine confidence was the normalized difference between a predefined threshold () and the sample’s score. Despite machines demonstrating superior accuracy (87.7%), they often exhibited lower confidence than humans. We counteracted this by scaling the machine’s confidence using a temperature parameter.
Fig.5 illustrates the trade-off between retaining human decisions post-machine intervention and the mean accuracy achieved, with Table4 (right part) breaking down the accuracy. At the intersection point in Fig. 5, humans make 56% of decisions (the rest are made by machines), achieving 84.5% accuracy, a performance boost of 16.4% over vanilla evaluation. This balance indicates that a synergistic approach between humans and AI can significantly enhance accuracy while ensuring that humans retain decision-making authority in the majority of cases.
Takeaways. Challenges enable human capabilities to detect deepfake audios, similar to machines, even on a significantly challenging dataset. Comparing human detection to machines reveals similarities and differences depending on the challenge type, suggesting a complementary relationship between human intuition and machine precision. When combined, machine tagging aided in assisting humans where their performance was lacking. Further collaboration between humans and machines resulted in an even greater boost in deepfake detection performance, while keeping humans in control.
6. Discussion
Usability. For a human-centered security system like Pitch, usability is paramount. Fig. 6 presents the willingness ratings of the 10 challenges as evaluated by 100 participants who contributed their data. For the caller-side, the median rating for performing all challenges was either Willing’ or Very Willing.’ As we vet all challenges for effectiveness, organizations can initially deploy our most usable challenges and incorporate others only when necessary.
We acknowledge that our proposed challenge-response framework incorporates manual components and requires callers to actively engage in a task before authentication. However, in terms of execution speed, it typically takes a caller mere 6 seconds to verbalize a sentence and complete the challenge.
On the defender’s (receiver’s) side, there is a desire for swift verification of callers before proceeding with the conversation. To accommodate this, we employed AI as a tagging system to aid the defender in making confident decisions. The design decision to maintain human control acknowledges the fallibility of AI and the nature of telephone communications, where a call, unlike spam emails, cannot be retrieved once disconnected.
Robustness against Countermeasures. An adaptive adversary may design countermeasures to counter our approach.
Firstly, given the public nature of challenges, adversaries can anticipate a challenge and prefabricate their responses, utilizing sophisticated editing tools to conceal detectable anomalies. However, our challenges feature randomizable elements, such as varying speech content and background noise, introduced at the call’s outset, to thwart pre-recordings.
Secondly, as deepfakes advance algorithmically, they may learn to bypass some of our current challenges. Anticipating this, defenders can employ a sequence of existing challenges, akin to current Captcha implementations. Defenders can also incorporate newly discovered challenges into Pitch to stay ahead of technological advancements in deepfake generation. Drawing from our own findings, we assert that there will always exist an audio challenge that deepfakes cannot faithfully generate, as the generative models do not simulate physics or encode human speech physiology.
Security-usability tradeoff: This tradeoff presents two perspectives: From the user’s perspective, we aim to minimize the number of challenges (usability). Conversely, from the defender’s perspective, we strive to make confident, accurate decisions (security).
Comparing detection performance (Fig. 2) with usability ratings (Fig. 4), we observe that certain challenges, such as cupping the mouth, whispering, or maintaining a static mouth, best balance this tradeoff. Even with a single challenge equivalent to verbalizing a single sentence ( seconds), we demonstrate high AUC (). Furthermore, if the defender requires greater certainty depending on the sensitivity of the phone interaction, they can request multiple challenges from the user to achieve that level of assurance.
Practical Design Recommendations. To ensure a practical implementation of the ‘Deepfake-Likely’ tagging system, we offer the following guidelines:
On the caller’s side:
Utilize single, succinct sentences for tasks.
Enhance compliance with high-quality instructional samples.
Avoid challenges that confuse both machines and humans, such as those involving human cognition tests or deception.
For playback, employ an automated sound for ease of use.
On the receiver’s side:
Employ two categories of tags: "Deepfake-Likely" and "Deepfake-Certainly," to indicate machine confidence.
Providing a rationale to back prediction has nominal benefits.
Limitations and Future Work. Our study has several limitations that provide directions for future research:
Comprehensive challenge evaluation: We did not assess all sub-categories in our challenge taxonomy, particularly those that might prove more relevant in a multilingual context, eg, roll R.
Testing against COTS deepfake generators: Pitch has not been tested against commercial off-the-shelf (COTS) deepfake generators like ElevenLabs.
Advanced machine learning models: Future work could utilize more sophisticated and recent foundation models to enhance cross-lingual machine detection accuracy.
Dataset diversity: Expanding our dataset to capture greater linguistic diversity, including speech variations like lisps, could improve the system’s robustness.
Inclusivity: Developing a more inclusive system that can calibrate detection thresholds for individuals with known speech impediments is an important addition.
7. Conclusion
This study presents a novel challenge-response approach to enhance both human and machine detection of interactive deepfake audio calls. We developed a comprehensive taxonomy of audio challenges, identifying 10 robust instances from 20 candidates, and provide a diverse open-source dataset. These challenges in the dataset effectively degraded deepfake quality, thereby improving machine detection capabilities and enabling humans to discern deepfakes, even against a challenging dataset baseline.
By fostering a synergistic relationship between humans and machines, we achieved a balance that capitalizes on human oversight and machine efficiency. This complementary approach has laid the foundation for an innovative tagging system aimed at pre-screening calls, marking a significant advancement in audio deepfake detection that operates independently of target identity.
Supplementary Material for Pitch
| General Sentences |
| The quick brown fox jumps over the lazy dog. |
| With measured words, she explained the strange rhythm of the universe. |
| Bright violets grow under the dense canopy, shading their vivid colors. |
| Jazz and classical music often feature complex harmonies and rhythms. |
| Six whimsical wizards make powerful, rhythmic chants deep in the forest. |
| The doctor’s prescription advised taking unique, holistic measures. |
| The swift cheetah is an elegant and extraordinary creature. |
| The rugged path wound around the steep hill to the quaint village. |
| Gazing at the azure sky in July, the dreamy poet wrote verses. |
| The glum monk thought deeply about life’s complex queries. |
| Questions |
| Does the quick brown fox jump over the lazy dog? |
| Can she explain the strange rhythm of the universe with measured words? |
| Do bright violets grow under the dense canopy, shading their vivid colors? |
| Do jazz and classical music often feature complex harmonies and rhythms? |
| Do six whimsical wizards make powerful, rhythmic chants deep in the forest? |
| Did the doctor’s prescription advise taking unique, holistic measures? |
| Is the swift cheetah an elegant and extraordinary creature? |
| Does the rugged path wind around the steep hill to the quaint village? |
| While gazing at the azure sky in July, did the dreamy poet write verses? |
| Did the glum monk think deeply about life’s complex queries? |
| Foreign Sentences |
| Finnish: ’En ymmarra mitaan mita tarkoitat’ |
| French: ’Le coeur a ses raisons que la raison ne connait point.’ |
| Spanish: ’El rapido zorro marron salta sobre el perro perezoso.’ |
| German: ’Fuchse lieben es, schnell uber die grauen Hunde zu springen.’ |
| Chinese: ’Feng yu zhi hou, bi jian cai hong.’ |
| Russian: ’Na dvore trava, na trave drova.’ |
| Arabic: ’Al-hayatu jamilatun ala al-raghmi min kulli shay’. |
| Hindi: ’Bandar bhagate hue chhat par chadh gaya.’ |
| Japanese: ’Sakura no hana ga haru no kaze ni mau.’ |
| Swahili: ’Simba anapenda kupumzika chini ya mti mkubwa.’ |
A.1. Data Collection Instructions
We requested each consenting participant to perform speech tasks (Table A.1) based on the following tasks.
0.
Speak the sentence normally.
Vocal Distortions
1.
Keep their mouth and lips static, and speak the sentence.
2.
Cup your mouth and Speak the sentence.
3.
Whisper the sentence.
Waveform Manipulation
4.
Speak the sentence loudly.
5.
Speak the sentence softly.
6.
Read out the sentence quickly.
7.
Read out the sentence slowly.
8.
Speak the sentence in a high pitch.
9.
Speak the sentence in a low pitch.
10.
Hold their nose and read the sentence.
Language and Articulation
11.
Speak a transliteration of a foreign sentence.
12.
Speak the sentence with a known accent, such as Russian or British.
13.
Sing the sentence.
Tone of Voice
14.
Speak the sentence with happiness or excitement.
15.
Speak the question.
Non-Lingual Noises
16.
Cough or Whistle.
17.
Speak the sentence while clapping.
Playback Challenges:
(done while playing audio on a smartphone) -
18.
Speak the sentence clearly amidst cross-talk; while playing recording of a matching sentence.
19.
Speak the sentence while playing instrumental music.
20.
Speak the sentence while playing lyrical music.
Desktop-users attempted all challenges, while smartphone users attempted everything except playback challenges (in our naive implementation, it requires another device to play the sound). Out of 20 listed tasks we found 10 that proved valuable against deepfakes.
| Challenge | DF1 | DF2 | VC1 | VC2 | ||
| No Challenge | 72.0 | 85.0 | 65.1 | 62.3 | 0.17 | 0.22 |
| Static Mouth | 64.2 | 88.5 | 92.5 | 86.9 | 0.19 | 0.84 |
| Cup mouth | 67.5 | 88.2 | 99.5 | 99.3 | -0.0 | 0.93 |
| Whisper | 84.8 | 90.1 | 90.7 | 95.9 | -0.12 | 0.97 |
| Speak softly | 70.5 | 82.6 | 75.5 | 79.3 | 0.09 | 0.79 |
| High Pitch | 73.8 | 81.9 | 90.3 | 82.8 | -0.01 | 0.90 |
| Foreign Words | 68.0 | 89.8 | 95.9 | 98.0 | 0.06 | 1.00 |
| Emotions | 78.9 | 82.7 | 81.0 | 80.1 | 0.22 | 0.90 |
| Cross-talk | 82.5 | 83.2 | 88.8 | 92.2 | -0.41 | 0.93 |
| Instr. Playback | 80.9 | 88.2 | 97.9 | 99.2 | -0.29 | 1.00 |
| Lyric Playback | 83.3 | 85.9 | 92.7 | 95.3 | -0.4 | 1.00 |
| Average | 75.1 | 86.0 | 88.2 | 88.3 | -0.05 | 0.86 |